OpenStudy (anonymous):

Time spent discussing material with students during office hours follow an exponential distribution with mean of 15 minutes. a) Find the probability that a student spends less than 10 minutes discussing material in office hours. b) Find the probability that a student spends more than 3 minutes discussing material during office hours. please and thanks been puzzling for me

OpenStudy (anonymous):

let me see if i can remember this correctly. i think we put \[\lambda =\frac{1}{15}\] and use \[P(X>s)=e^{-\lambda s}\] and \[P(X<s)=1-e^{\lambda s}\] does that look familiar?

OpenStudy (anonymous):

sorry last one should be \[P(X<s)=1-e^{-\lambda s}\]

OpenStudy (anonymous):

so first one would be \[P(X<10)=1-e^{-\frac{10}{15}}=1-e^{-\frac{2}{3}}=0.48658288\] rounded

OpenStudy (anonymous):

similarly \[P(X>3)=e^{-\frac{3}{15}}=e^{\frac{-1}{5}}=0.8187\]

OpenStudy (anonymous):

thanks , yes that is the correct value (1/15) i do believe your on the right track

OpenStudy (anonymous):

just looked it up and they are right unless i really missed my guess

OpenStudy (anonymous):

thanks your correct, its concistent with my notes...you wouldnt possibly know how to the previous one i posted would you?

OpenStudy (anonymous):

i will look